OpenStudy (smarty):

[2 + 3^2 – 3] ÷ 2 • 4

OpenStudy (anonymous):

20

OpenStudy (smarty):

The choices are 16 and 1.

OpenStudy (anonymous):

Sixteen. The sum in the brackets is eight, divided by two is four times four is sixteen.

OpenStudy (smarty):

@AnimalAin What about 8/8?

OpenStudy (anonymous):

The last factor (the four) is multiplied, not divided. It will be in the numerator, not the denominator. If the author wanted it to be in the denominator, it would have parentheses around the two and four.

OpenStudy (anonymous):

\[[2 + 3^2 – 3] ÷ 2 • 4=\frac{8*4}{2}=16\]

OpenStudy (smarty):

ok thanks!

OpenStudy (anonymous):

No sweat. Do math every day.